# Quantum Gravity

### v1

@@ Abstract
We investigate the local limits of various classes of unitary, nonlocal quantum field theories. While it is easy to build nonlocal models with well-behaved asymptotics in Euclidean space, the Minkowskian correlation functions typically exhibit singular behaviors. We introduce “asymptotically local” quantum field theory (AL-QFT) as the class that encompasses unitary, nonlocal theories with well-defined local limits in Minkowski spacetime. The target models cannot propagate ghosts, but are allowed to contain purely virtual particles (PVPs). In the bubble diagram, the nonlocal deformation generates PVPs straightforwardly. In the triangle diagram, it does so possibly up to multi-threshold corrections, which may be adjusted by tuning the deformation itself. We build examples of AL-QFTs, including a deformation of quantum gravity with purely virtual particles. AL-QFT can serve various purposes, such as suggesting innovative approaches to off-shell physics, providing an alternative formulation for theories with PVPs, or smoothing out nonanalytic behaviors. We discuss its inherent arbitrariness and the implications for renormalizability.
@@ 1 Introduction
The space $\mathcal{M}_{\text{NL}}$ of nonlocal theories (NL-QFTs) is huge. The space $\mathcal{M}_{\text{L}}$ of local theories (L-QFTs) may be understood as a subspace of its boundary $\partial\mathcal{M}_{\text{NL}}$. It may be helpful to organize $\mathcal{M}_{\text{NL}}$ into subspaces that have interesting properties, with special focus on the neighborhood $\mathcal{U}_{\text{NL/L}}$ of $\mathcal{M}_{\text{L}}$.
@@
Various classes of nonlocal theories have been studied in the literature, starting from Efimov [1]. Krasnikov [2] had the idea to remove the ghosts of local theories from the spectrum by means of nonlocal deformations. Specifically, the ghost poles of the free propagators are canceled by “form factors”, that is to say, entire functions with no zeros that appear instead of the unwanted poles. In so doing, unitarity is ensured [3, 4].
@@
If the form factors are generic, they originate nonrenormalizable gauge and gravity interactions. Concentrating on gravity, Kuz’min [5] showed that renormalizability can be obtained by choosing entire functions that tend to polynomials at high energies. The parent and deformed theories are super-renormalizable. Precisely, the nonlocal propagators and vertices tend sufficiently fast to those of a higher-derivative (HD) theory in the ultraviolet limit [6]. Later Tomboulis [7] revived this idea and applied it to gauge theories. More recently, Modesto and others [8] elaborated it to a greater extent and built a variety of finite models. Properties and implications of more general form factors were studied in ref.s [9].
We take the class of super-renormalizable nonlocal theories just mentioned, which we call $\mathcal{M}_{\text{NL}}^{0}$, as the starting point of our investigation. Assuming it makes sense to discuss a “local limit” in $\mathcal{M}_{\text{NL}}^{0}$, we expect it to recover the parent HD-QFT. However, the relationship between nonlocal and local theories has not been investigated so far to the extent we aim to explore in this paper.
We assume that the manifolds $\mathcal{M}_{\text{NL}}$ and $\mathcal{M}_{\text{L}}$ are restricted to contain unitary theories only. The first step is to define what we mean by “local limit” in $\mathcal{M}_{\text{NL}}$. We consider a family of theories in $\mathcal{M}_{\text{NL}}$ parametrized by some extra variable $\lambda$. We arrange the $\lambda$ dependence so that the limit $\lambda\rightarrow\infty$ returns a local theory in Euclidean space, where the problem is much simpler to handle. Then we inquire what happens when $\lambda$ tends to infinity in Minkowski spacetime. In a class of treatable theories inspired by the ones of $\mathcal{M}_{\text{NL}}^{0}$ (but not quite those), we find severe singularities inside the correlation functions, such as integrals $\simeq\int\mathrm{d}^{4}p/|p^{2}-m^{2}|$. We are unable, at this very moment, to say whether the nonlocal models of $\mathcal{M}_{\text{NL}}^{0}$ admit local limits in Minkowski spacetime or not.
Still, the results we find suggest that the usual assumptions behind NL-QFT are too restrictive. We expect that if we adjust, or enlarge, the manifold of nonlocal models, we can include those that tend to local quantum field theories in Minkowski spacetime when $\lambda$ tends to infinity.
A question that may help us identify the right extension is: which local models can be reached and which ones are out of reach? Since we want $\mathcal{M}_{\text{NL}}$ to be unitary, the local models that contain ghosts are unreachable. Then, the natural candidates for the intersection $\partial\mathcal{M}_{\text{NL}}\cap\mathcal{M}_{\text{L}}$ are the local theories with purely virtual particles [10] (i.e., particles that are never on the mass shell), alongside physical particles.
@@
We need to understand the relation between the local limit $\lambda\rightarrow\infty$, and the continuation E $\rightarrow$ M from Euclidean space (E) to Minkowski spacetime (M). Do these operations commute? If we take the local limit in E and then continue from E to M, do we find the local limit in M?
@@
Unless we assume that it is analytic, the continuation E $\rightarrow$ M is not unique, even if we restrict it to $\mathcal{M}_{\text{L}}$. However, the analytic continuation generically leads to models with ghosts (fields with kinetic terms multiplied by the wrong sign). This occurs in the cases of higher-derivative theories, for example. Since those models are unreachable from $\mathcal{M}_{\text{NL}}$, we conclude that the continuation E $\rightarrow$ M, which is analytic for $\lambda<\infty$, cannot stay analytic after the local limit $\lambda\rightarrow\infty$. There remains the possibility that the target local limits are theories involving purely virtual particles (PVPs). Indeed, a nonanalytic continuation E $\rightarrow$ M is one way to formulate PVPs [11].
@@
Two other ways to introduce purely virtual particles are available in the literature: one uses diagrammatic spectral optical identities [12] and the other one is based on special non-time ordered correlation functions [10]. A fourth way is the one emerging from this paper. Generically speaking, we can say that PVPs are defined by tweaking the usual diagrammatics so as to eradicate the ghosts, while remaining in the realm of local theories.
@@
Taking advantage of the PVP concept, it is possible to build a local theory of quantum gravity [13] that is unitary and renormalizable at the same time. Its main prediction is a constrained window ($10004/10000\lesssim r\lesssim 3/$ 1000) for the value of the tensor-to-scalar ratio $r$ [14] of primordial fluctuations, which essentially confirms the prediction of the Starobinsky $R+R^{2}$ model [15] ($r\simeq 3/1000$) within less than an order of magnitude. Other predictions can be derived as well (the running of scalar and tensor tilts, higher-order corrections, etc. [16]), but require much more effort to be tested, in the realm of current or planned observations [17]. In this paper we provide a nonlocal deformation of quantum gravity with PVPs, which returns it back in the local limit.
Since we know what to expect (i.e., local limits containing physical particles and PVPs), we can easily identify how to relax the defining assumptions of NL-QFT to enlarge the manifold $\mathcal{M}_{\text{NL}}$ appropriately. It turns out that the form factors attached to the “propagators of non propagating fields”, which eliminate the unwanted ghost poles, remain entire functions, but are allowed to have zeros. This generalization does not cause particular problems, since the associated fields have to be integrated out anyway, sooner or later. The Lagrangian is singular, strictly speaking, but the action is still regular.
We show that the so enlarged manifold $\mathcal{M}_{\text{NL}}$ does include models that have regular local limits in Minkowski spacetime, and those limits are indeed theories that contain purely virtual particles in addition to physical particles (PVP-QFTs). We call the models of the subspace $\mathcal{U}_{\text{NL/L}}\subset\mathcal{M}_{\text{NL}}$ “asymptotically local” quantum field theories (AL-QFTs), and denote their space by $\mathcal{M}_{\text{AL}}$ henceforth. AL-QFT can be used as an alternative formulation of theories with PVPs, or as an approximation that smooths out their typical nonanalytic behaviors.
@@
Besides the tree level, where the local limit is straightforward, we study the local limits of the bubble and triangle diagrams. In the bubble diagram the nonlocal deformation “already knows”, so to speak, how to generate PVPs in the limit, with no need of ad hoc adjustments. In the cases of triangle and more complicated diagrams, the local limit is “PVP ready” in the null and single-threshold sectors. The multi-threshold sectors, instead, are more difficult to handle. Possibly, the arbitrariness of the deformation must be invoked to fine-tune the limit appropriately.
@@
The arbitrariness of AL-QFT is its weak point, which must be addressed. Nonlocal quantum field theory lacks a fundamental principle to select the nonpolynomial functions it is built upon, remove its inherent arbitrariness and identify the unique theory that describes nature. At the same time, this weakness is predicated on the very assumption that AL-QFT is ambitious to that point. This is not what we are claiming here: we are not proposing AL-QFT as a framework for fundamental theories of the universe. We just claim that AL-QFT is interesting per se, broadens our knowledge of QFT, and provides useful tools in support of local QFT.
To clarify this last point, let us recall another instance where a comparable degree of arbitrariness enters local QFT (without jeopardizing it as a framework for fundamental interactions): we are talking about off-the-mass shell physics [18]. In that case, the extra parameters are not rooted in the fundamental interactions, but describe the surrounding environment where the phenomenon is observed. We can say that they parametrize the quantum/classical interplay between the phenomenon, the observer, and the experimental apparatus. There is probably a map relating the arbitrariness of AL-QFT to the arbitrariness of off-shell physics. Yet, elucidating that map is beyond the scope of this paper. Here we just stress that the arbitrariness turned on by the deformation PVP-QFT $\rightarrow$ AL-QFT is less surprising when it is placed side by side with other forms of arbitrariness we are more accustomed to.
Summarizing, the lack of uniqueness of AL-QFT is not an issue, within our approach. AL-QFT is a tool to enlarge the set of quantum field theories we can treat, move beyond the common frameworks, possibly describe off-the-mass-shell physics in an alternative way, or propose new formulations of PVPs, and address peculiar aspects of PVPs themselves (such as the violation of microcausality [19] and the so called “peak uncertainty” [20]).
@@
Another way to settle the non uniqueness problem of AL-QFT is as follows. We have said that nonlocal theories lack a selection criterion for the form factors, a sort of “minimum principle” in $\mathcal{M}_{\text{AL}}$. Yet, the requirement that they admit local limits provides a form of control, and possibly the missing principle itself. In this view, the “minima”, i.e., the candidate theories to describe nature from AL-QFT, are nothing but the local limits themselves [^1] Strictly speaking, they are not contained in the manifold $\mathcal{M}_{\text{AL}}$ of asymptotically local theories, but in its border $\partial\mathcal{M}_{\text{AL}}$..
The paper is organized as follows. In section 2 we recall basic features of the nonlocal theories considered in the literature and outline how we generalize them. In section 3 we study the local limit in Euclidean space. In section 4 we describe the problem posed by the local limit in Minkowski spacetime. In section 5 we study the bubble and triangle diagrams. In section 6 we formulate the asymptotically local deformation of quantum gravity with purely virtual particles. Section 7 contains the conclusions. In the appendix we recall the calculation of the bubble diagram with PVPs.
@@
When necessary, the dimensional regularization [21] is used for the explicit calculations, where $D=4-\varepsilon$ denotes the continued dimension around dimension 4.
2 Nonlocal theories
Nonlocal theories must be defined in Euclidean space and later analytically continued to Minkowski spacetime. We consider models with Euclidean Lagrangians of the form
@@ 2 Nonlocal theories
in momentum space, where $\phi$ denotes the fields, $Q(P)$ is a certain function of a polynomial $P$ of the Euclidean momentum $p$, and $\tau=\pm 1$ (to describe nonlocal deformations of both physical particles and ghosts). Typically, we take
@@
where $m$ is the “mass” in the local limit.
We require that $Q(P)$ tends to $P$ in the local limit defined below and $1/Q(P)$ is entire in the complex $P$ plane. The idea is that the field $\phi$ is itself purely virtual [^2] We could call it “nonlocal purely virtual particle” (NL-PVP)., since $Q(P)$ has no zeros. This makes it a natural candidate to become a PVP in the local limit.
The main difference with respect to the assumptions commonly adopted in the literature about nonlocal theories [5, 7, 8] is that the propagator $\tau/Q(P)$ is not required to never vanish in the complex $P$ plane. Specifically, we allow $Q(P(p))$ to have singularities proportional to $(-(p_{\text{M}}-k_{\text{M}})^{2}+M^{2})^{-1}$, for Minkowskian momenta $p_{\text{M}}$, $k_{\text{M}}$, and real “masses” $M$. The Minkowskian action is then defined by means of the Cauchy principal value.
@@
The local limit is defined by rescaling $P$ and $Q$ by $\lambda$ and $\lambda^{-1}$, respectively, where $\lambda$ is a positive factor, and then letting $\lambda$ tend to infinity. We assume that $Q$ tends to $P$ on the real axis:
@@
where $h(P)$ is the entire function defined in formula (2.6) below, taken from the current literature on nonlocal theories [2, 5, 7, 8]. The second option allows us to build the asymptotically local theories of $\mathcal{M}_{\text{AL}}$. With choices like (2.5), the property (2.4) extends to a double cone $\mathcal{C}$ around the real axis.
2.1 Approximating the absolute value
$h(z)\equiv\exp\left(\frac{1}{2}\int_{0}^{z^{2}}\frac{1-\mathrm{e}^{-w}}{w}\mathrm{d}w-\frac{\gamma_{\text{E}}}{2}\right)=\exp\left(\frac{1}{2}\ln z^{2}+\frac{1}{2}\Gamma(0,z^{2})\right).$ (2.6)
Precisely, in the double cone $\mathcal{C}=\{z:-\pi/4<$ arg[$z$] $<\pi/4$ or $3\pi/4<$ arg[$z$] $<5\pi/4\}$, we have [22]
$h(z)=\sqrt{z^{2}}\left[1+\frac{\mathrm{e}^{-z^{2}}}{2z^{2}}\left(1+\mathcal{O}\left(\frac{\mathrm{e}^{-z^{2}}}{z^{2}}\right)\right)\left(1+\mathcal{O}\left(\frac{1}{z^{2}}\right)\right)\right],$ (2.7)
so
$\lim_{\lambda\rightarrow+\infty}\frac{h(\lambda z)}{\lambda}=\sqrt{z^{2}}\text{\quad for }z\in\mathcal{C},\qquad\lim_{\lambda\rightarrow+\infty}\frac{h(\lambda x)}{\lambda}=|x|\quad\text{\ for }x\in\mathbb{R}.$ (2.8)
It is important to stress that $h(z)$ has no zeros. Its reciprocal $1/h(z)$ is going to be useful to build the “propagators” that tend to purely virtual particles in the local limit.
The function $h(z)$ is even,
$h(z)=h(-z),$ (2.9)
and satisfies
$h^{\ast}(z)=h(z^{\ast}).$ (2.10)
Moreover, on the real axis it is positive and bounded from below:
$h(x)\geqslant h(0)|x|,\qquad x\in\mathbb{R}\text{,}$ (2.11)
where $h(0)=\mathrm{e}^{-\gamma_{\text{E}}/2}=0.7493...$ The plot of fig. 1 makes this bound transparent.
Finally, (2.8) gives
$\lim_{\lambda\rightarrow+\infty}\frac{h^{2}(\lambda z)}{\lambda^{2}z}=z\quad\text{for }z\in\mathcal{C}.$ (2.12)
Figure 1: Plot of $h(x)$ for real $x$
2.2 Approximating the sign function
The sign function sgn$(x)$ can be approximated on the complex plane by the entire function
$\sigma(z)=\frac{z}{h(z)},$ (2.13)
$\lim_{\lambda\rightarrow+\infty}\sigma(\lambda z)=\text{sgn}(\text{Re}[z])\quad\text{for }z\in\mathcal{C},\qquad\lim_{\lambda\rightarrow+\infty}\sigma(\lambda x)=\text{sgn}(x)\quad\text{for }x\in\mathbb{R}.$ (2.14)
Note that $\sigma(z)$ is odd in the complex plane:
$\sigma(z)=-\sigma(-z).$ (2.15)
Moreover, (2.11) gives
$|\sigma(x)|\leqslant\frac{1}{h(0)},\qquad x\in\mathbb{R}\text{.}$ (2.16)
2.3 Approximating the principal value
We can use the functions $\sigma$ and $h$ to approximate the Cauchy principal value $\mathcal{P}$ on the real axis. Specifically, we prove that
$\lim_{\lambda\rightarrow+\infty}\frac{\lambda\sigma(\lambda x)}{h(\lambda x)}=\mathcal{P}\frac{1}{x},\qquad x\in\mathbb{R},$ (2.17)
in the sense of distributions.
First observe that
$\lim_{\lambda\rightarrow+\infty}\frac{\lambda\sigma(\lambda x)}{h(\lambda x)}=\frac{1}{x}$ (2.18)
for every real $x\neq 0$. This result follows from (2.8) and (2.14).
If $\varphi(x)$ denotes a real test function, consider the integral
$\mathcal{I}\equiv\lim_{\lambda\rightarrow+\infty}\int_{-\infty}^{+\infty}\mathrm{d}x\frac{\lambda\sigma(\lambda x)\varphi(x)}{h(\lambda x)},$ (2.19)
which can also be written as
thanks to (2.9) and (2.15). The modulus of the $\mathcal{I}$ integrand is bounded above by a $\lambda$-independent function that is integrable on $\mathbb{R}$. Indeed, the inequalities (2.11) and (2.16) give
$\left|\frac{\lambda\sigma(\lambda x)(\varphi(x)-\varphi(-x))}{2h(\lambda x)}\right|\leqslant\frac{1}{2h^{2}(0)}\left|\frac{\varphi(x)-\varphi(-x)}{x}\right|,$ (2.20)
for real $x$. Then the dominated convergence theorem allows us to exchange the limit with the integral. Given that (2.18) holds almost everywhere, we obtain
$\mathcal{I}=\int_{-\infty}^{+\infty}\mathrm{d}x\frac{\varphi(x)-\varphi(-x)}{2x}=\mathcal{P}\int_{-\infty}^{+\infty}\frac{\mathrm{d}x}{x}\varphi(x),$ (2.21)
as we wished to show.
2.4 Loop integrals in nonlocal theories
The propagator of the nonlocal Euclidean theory (2.1) is

### v3

@@ Abstract
We investigate the relationship between nonlocal and local quantum field theories, and search for a viable notion of “local limit” to relate the unitary models. In Euclidean space it is relatively easy to have nonlocal theories with well-behaved local limits. In Minkowski spacetime, instead, singular behaviors are generically expected. Relaxing some assumptions on the “form factors” considered in the literature, we identify a class of models that have regular local limits in Minkowski spacetime. We call the models “asymptotically local” quantum field theories (AL-QFTs) and show that their limits are theories with physical and purely virtual particles (PVPs). In the bubble diagram, the nonlocal deformation generates PVPs straightforwardly. In the triangle diagram, it does so possibly up to multi-threshold corrections, which may be adjusted by tuning the deformation itself. We also build an asymptotically local deformation of quantum gravity with purely virtual particles. AL-QFT can serve various purposes, such as suggesting innovative approaches to off-shell physics, providing an alternative formulation for theories with PVPs, or smoothing out nonanalytic behaviors. We discuss its inherent arbitrariness and the implications for renormalizability.
@@ 1 Introduction
In this paper we consider various possibilities in this regard. The main goal is to work out a practically viable notion of “local limit” to relate local and nonlocal theories with physically desirable properties, such as renormalizability and unitarity. Besides deepening our knowledge on quantum fields, this investigation may have interesting physical applications. Nonlocal deformations of the standard model, for example, may be suitable to describe off-shell physics. To our knowledge, the relationship between nonlocal and local theories has not been investigated, so far, to the extent we aim to explore in this paper.
@@
Various classes of nonlocal theories have been studied in the literature, starting from Efimov [1]. Krasnikov [2] had the idea to remove the ghosts of local theories from the spectrum by means of nonlocal deformations. Specifically, the unphysical poles of the free propagators are replaced by certain form factors (entire functions with no zeros) to ensure unitarity [3, 4].
@@
If the form factors are generic, they originate nonrenormalizable gauge and gravity interactions. Concentrating on gravity, Kuz’min [5] showed that renormalizability can be obtained by choosing entire functions that tend to polynomials at high energies. The parent local and the deformed nonlocal theories are super-renormalizable. Precisely, the propagators and vertices of the latter tend sufficiently fast to those of a higher-derivative (HD) theory in the ultraviolet limit [6].
Later, Tomboulis [7] revived this idea and applied it to gauge theories. More recently, Modesto and others [8] elaborated it to a greater extent and built a variety of finite models. Properties and implications of more general form factors were studied in ref.s [9].
The class of super-renormalizable nonlocal theories just mentioned provides an interesting arena for testing candidate notions of local limit. Naively, we may expect that the limit is just the higher-derivative local theory (HD-QFT) that describes the ultraviolet behavior. It turns out that the relationship between locality and nonlocality is more subtle than that.
Let $\mathcal{M}_{\text{NL}}$ and $\mathcal{M}_{\text{L}}$ denote the manifolds of nonlocal and local unitary theories (NL-QFTs and L-QFTs, from now on), respectively. We look for families of models in $\mathcal{M}_{\text{NL}}$, parametrized by some variable $\lambda>0$, that become local when $\lambda\rightarrow\infty$. This way, the space $\mathcal{M}_{\text{L}}$ may be understood as a subspace of the boundary $\partial\mathcal{M}_{\text{NL}}$.
It is relatively easy to provide a meaningful notion of local limit in Euclidean space. It is more difficult to extend the concept to Minkowski spacetime. For example, in a class of treatable theories inspired by the ones of the literature mentioned above, we find that the limit generates severe singularities inside the correlation functions, such as integrals $\simeq\int\mathrm{d}^{4}p/|p^{2}-m^{2}|$. We are unable, at this very moment, to say whether the nonlocal models of refs. [5, 7, 8] admit proper local limits in Minkowski spacetime or not.
This leads us to search for new, ad hoc nonlocal models that do have meaningful local limits in Minkowski spacetime when $\lambda$ tends to infinity. We find that the form factors attached to the “propagators of nonpropagating fields” – those which eliminate the unwanted poles – must have some zeros. Apart from that, they are still entire functions. This generalization does not do too much harm, since the associated fields can be integrated out with no loss. The Lagrangian is singular, strictly speaking, but the action is regular.
@@
Another direction to pursue to shed light on the matter is the relation between the local limit $\lambda\rightarrow\infty$ and the continuation E $\rightarrow$ M from Euclidean space (E) to Minkowski spacetime (M). Do these operations commute? If we take the local limit in E and then continue from E to M, do we find the local limit in M?
@@
The continuation E $\rightarrow$ M is not unique. In $\mathcal{M}_{\text{L}}$, for example, the analytic continuation of higher-derivative theories gives ghosts (fields with kinetic terms multiplied by the wrong sign), while the “average continuation” [10] gives theories with purely virtual particles (PVPs), i.e., particles that are always off shell. By unitarity, the ghosts are unreachable from $\mathcal{M}_{\text{NL}}$, so the continuation E $\rightarrow$ M, which is analytic for $\lambda<\infty$, cannot stay analytic when $\lambda\rightarrow\infty$. The possibility that the target local limits contain PVPs is compatible with this.
Summarizing, in this paper we build “asymptotically local” quantum field theories (AL-QFTs) – models that have regular local limits in Minkowski spacetime. Such limits contain purely virtual particles in addition to physical particles (PVP-QFTs).
@@
We recall two additional methods of introducing purely virtual particles, besides the one already mentioned: one uses diagrammatic spectral optical identities [11] and the other one is based on non-time ordered correlation functions of a special type [12]. Generically speaking, PVPs are defined by tweaking the diagrammatics so as to eradicate the ghosts, while remaining in the realm of local theories. That is to say, the PVP and Krasnikov approaches share the same goals, in local and nonlocal settings, respectively. We can also view AL-QFT as a fourth formulation of theories with PVPs.
@@
We also mention that, taking advantage of the PVP concept, it is possible to build a local theory of quantum gravity [13] that is unitary and renormalizable at the same time. Its main prediction is a constrained window ($4/10000\lesssim r\lesssim 3/1000$) for the value of the tensor-to-scalar ratio $r$ [14] of primordial fluctuations, which essentially confirms the prediction of the Starobinsky $R+R^{2}$ model [15] ($r\simeq 3/1000$) within less than an order of magnitude. Other predictions can be derived as well (the running of scalar and tensor tilts, higher-order corrections, etc. [16]), but require more effort to be tested, in the realm of current or planned observations [17]. Among the other things, in this paper we provide a nonlocal deformation of quantum gravity with PVPs, which returns the theory of [13] in the local limit.
@@
The local limit is straightforward at the tree level, where most of its intriguing aspects are not visible. We push the investigation to the simplest nontrivial diagrams, that it to say, the bubble and the triangle. We show that in the case of the bubble diagram the nonlocal deformation “already knows”, so to speak, how to generate PVPs, with no need of ad hoc adjustments. In the cases of triangle and more complicated diagrams, the local limit is “PVP ready” in the null and single-threshold sectors. The multi-threshold sectors, instead, are more difficult to handle. Possibly, the arbitrariness of the deformation must be invoked to fine-tune the limit appropriately.
@@
The weakness of nonlocal quantum field theory is that it lacks a fundamental principle to select the nonpolynomial functions it is built upon, remove its inherent arbitrariness and identify a unique candidate theory of nature. However, we do not propose AL-QFT as a framework for fundamental theories. We just claim that AL-QFT is interesting per se, broadens our knowledge of QFT, and provides useful tools in support of local QFT. For example, as an alternative formulation of PVPs it may allow us to address peculiar aspects of these “non particles” (such as the violation of microcausality [18] and the so called “peak uncertainty” [19]).
Another instance where a comparable degree of arbitrariness plagues (local) QFT (without jeopardizing it as a framework for fundamental interactions) is off-the-mass shell physics [20]. There, the extra parameters describe the environment where the phenomenon under observation is placed. AL-QFT may provide an alternative way to describe these aspects. In this respect, the lack of uniqueness of AL-QFT is not an issue.
@@
Another way to settle the non uniqueness of AL-QFT is as follows. We have remarked that nonlocal theories lack a selection principle for the form factors. The local limit might serve as one. Then, the candidate theories to describe nature from AL-QFT are nothing but their local limits [^1] Strictly speaking, they are not contained in the manifold $\mathcal{M}_{\text{AL}}$ of asymptotically local theories, but in its border $\partial\mathcal{M}_{\text{AL}}$..
The paper is organized as follows.
  1. 1.
    In section 2 we recall the basic features of the nonlocal theories considered in the literature and explain how we are going to generalize them to have well-defined local limits in Minkowski spacetime.
  2. 2.
    In section 3 we study the local limit in Euclidean space, which is relatively straightforward.
  3. 3.
    In section 4 we describe the local limit in Minkowski spacetime, including explicit calculations for the bubble and triangle diagrams. We compare the behaviors of the nonlocal theories of the liturature with the properties of the new models built in section 2.
  4. 4.
    In section 5 we formulate an asymptotically local deformation of quantum gravity with purely virtual particles along the lines of section 2.
@@
Section 6 contains the conclusions. In appendix A we build the key functions that are needed for the form factors used in the paper. In appendix B we recall the calculation of the bubble diagram with PVPs. When necessary, the dimensional regularization [21] is used for explicit calculations, where $D=4-\varepsilon$ denotes the continued dimension around dimension 4.
2 The nonlocal theories we consider here
We want to study the local limits of nonlocal theories. There is no guarantee that such a limit exists, in general, so we may have to construct ad hoc models that admit one.
In this section we lay out the basic features of the classes of theories under consideration. We include those mostly studied in the current literature, as well as extensions that have well-behaved local limits.
We first work in Euclidean space and later switch to Minkowski spacetime by means of the analytic continuation. The models we focus on have Euclidean Lagrangians of the form
@@ 2 The nonlocal theories we consider here
in momentum space, where $\phi$ denotes the fields, $Q(P)$ is a certain function of a polynomial $P$ of the Euclidean momentum $p$, and $\tau=\pm 1$ is there so that we can describe nonlocal deformations of both physical particles and ghosts.
We take
@@
where $m$ is the “mass” in the local limit. Polynomials of higher-derivative theories, such as
$P_{\text{HD}}(p)=(p^{2}+m^{2})^{2}+M^{4},$ (2.3)
are also interesting, but will not be treated in detail here.
For a while, we concentrate on the propagator and study the local limit there.  We assume that the vertices are local or their limits do not invalidate the derivations we make. We discuss the vertices in detail in subsection 5.4.
In this context, the key quantity that encodes the nonlocal model is the function $Q(P)$. If we want a degree of freedom from $\phi$, we know how to obtain it: $Q=P$. The challenge is to not have a degree of freedom from $\phi$. That is to say, $\phi$ should be purely virtual [^2] We could call it “nonlocal purely virtual particle” (NL-PVP).. To achieve this, we require that $\tau/Q(P)$ is entire in the complex $P$ plane, so that $Q(P)$ has no zeros. In addition, we require that $Q(P)$ tends to $P$ in the local limit defined below. This makes $\phi$ a natural candidate to become an ordinary PVP in that limit.
The main difference with respect to the assumptions commonly adopted in the literature [5, 7, 8] is that the propagator $\tau/Q(P)$ is not required to never vanish in the complex $P$ plane. Specifically, we allow $Q(P(p))$ to have a singularity proportional to $(-p_{\text{M}}^{2}+M^{2})^{-1}$, where $p_{\text{M}}$ is the Minkowskian momentum and $M$ is some mass. The Minkowskian action is then defined by means of the Cauchy principal value.
@@
2.1 Local limit
To begin with, we define the local limit as follows: we rescale $P$ and $Q$ by $\lambda$ and $\lambda^{-1}$, respectively, where $\lambda$ is a positive factor, and then let $\lambda$ tend to infinity. We assume that $Q$ tends to $P$ on the real axis:
@@ 2.1 Local limit
where $h(P)$ is the entire function defined in formula (A.1), taken from the current literature on nonlocal theories [5, 7, 8].
In appendix A we show that $h(P)$ tends to the absolute value, $P/h(P)$ tends to the sign function and $P/h^{2}(P)$ tends to the Cauchy principal value in the local limit. The second option of (2.6) is the one that allows us to build the asymptotically local models.
With choices like (2.6), the property (2.5) extends to a double cone $\mathcal{C}$ around the real axis (see appendix A).
2.2 Propagator
The propagator of the Euclidean theories (2.1) is